Engineering Logistics

An electronics company has two contract manufacturers in Asia: Foxconn assembles its tablets and smart phones and Flextronics assembles its laptops. Monthly demand for tablets and smartphones is 15,000 units, whereas that for laptops is 6,000. Tablets cost the company $150, laptops cost $500, and the company has an annual holding cost of 20 percent. Currently the company has to place separate orders with Foxconn and Flextronics and receives separate shipments. The fixed cost of each shipment is $20,000.

  1. What is the optimal order size and order frequency with each of Foxconn and Flextronics? Show your computations clearly.
  2. The company is thinking of combining all assembly with the same contract manufacturer. This will allow for a single shipment of all products from Asia. If the fixed cost of each shipment remains $10,000, what is the optimal order frequency and order size from the combined orders? How much reduction in cycle inventory can the company expect as a result of combining orders and shipments? Show your computations clearly.
Answer #1

a.

Foxconn:

Monthly Demand for tablets = 15000

Annual Demand, D = 15000*12 = 180000

Cost of tablets = 150

Annual Holding Cost rate = 20%

Annual Holding Cost, H = 150*20% = 30

Fixed cost of shipment, S = 20000

Optimal Order Size , Q= \sqrt{} (2DS/H) = \sqrt{} (2*180000*20000/30) = 15491.93

Order Frequency = D/Q = 180000/15491.93 = 11.62

Foxtronics:

Monthly Demand for laptops = 6000

Annual Demand, D = 6000*12 = 72000

Cost of laptops = 500

Annual Holding Cost rate = 20%

Annual Holding Cost, H = 500*20% = 100

Fixed cost of shipment, S = 20000

Optimal Order Size , Q= \sqrt{} (2DS/H) = \sqrt{} (2*72000*20000/100) = 5366.56

Order Frequency = D/Q = 72000/5366.56 = 13.42

b. Combined Shipment:

Monthly Demand = 15000+6000 = 21000

Annual Demand, D = 21000*12 = 252000

Monthly Inventory Value for Tablets = Monthly demand*cost = 15000*150 = 2250000

Monthly Inventory Value of laptops = 6000*500 = 3000000

Total inventory value of the combined shipment = 2250000+3000000 = 5250000

Average per unit inventory value of the combined shipment = 5250000/21000 = 250

Holding cost per unit , H= 250*20% = 50

Fixed cost per order, S = 10000

Optimal Order Size, Q =  \sqrt{}(2DS/H) = \sqrt{} (2*252000*10000/50) = 10039.92

Order Frequency = D/Q = 252000/10039.92 = 25.09

Cycle inventory = Q/2

Cycle inventory of tablets = 15491.93/2 = 7745.96

Cycle inventory of laptops = 5366.56/2 = 2683.28

The total cycle inventory if ordered separately = 7745.96+2683.28 = 10429.25

Cycle inventory of combined shipment = 10039.92/2 = 5019.96

The reduction of cycle inventory the company can expect as a result of combined shipments = 10429.25-5019.96 = 5409.28